OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN! Algebra Find the number that should be added to the expression to make it a perfect square

OpenStudy (anonymous):

@jim_thompson5910 

OpenStudy (anonymous):

so u know that (x+y)^2 = x^2+2xy+y^2 ?

jimthompson5910 (jim_thompson5910):

take half of the x coefficient then square that result to get your answer

OpenStudy (anonymous):

so half means divide by 2? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes to cut something in half means to divide by 2 eg 1/2 of 10 = 10 divided by 2 = 5

OpenStudy (anonymous):

okay so half of 2 is one?

jimthompson5910 (jim_thompson5910):

the x coefficient is -6, not 2

jimthompson5910 (jim_thompson5910):